OpenStudy (kainui):

Medal to the first person to answer correctly! Evaluate this:

OpenStudy (kainui):

\[\LARGE \Re({ e^z})\]

OpenStudy (anonymous):

whats r and whats z

OpenStudy (kainui):

R means evaluate the real part, z is a complex number.

OpenStudy (kainui):

Whoops I messed up, oh well. I meant to write \[\LARGE \Re({ e^{iz}})\] But it's too late now lol

OpenStudy (freckles):

so R(a+bi)=a? that is the definition of that function?

OpenStudy (anonymous):

I tried this the other day and I got \[ e^{-\Im(z)} \cos(\Re(z)) \]

OpenStudy (anonymous):

But I don't think it was right.

OpenStudy (freckles):

where a and b are real

OpenStudy (kainui):

Not quite, the real and imaginary parts exist separately as a linear combination as a+bi. a is the real and b is the imaginary.

OpenStudy (freckles):

what does evaluates the real part mean

OpenStudy (kainui):

So if I ask for the real part of z^2 then I would do this \[\Large z^2 = (a+bi)^2 = a^2+b^2 -i2ab\] So we write \[\Large \Im(z^2) = -2ab \\ \Large \Re(z^2) = a^2+b^2\]

OpenStudy (freckles):

so R(a+bi)=a then?

OpenStudy (kainui):

Yep

OpenStudy (freckles):

ok i got confused for a second when you said not quite

OpenStudy (kainui):

Oh I was answering wio, sorry for the confusion

OpenStudy (freckles):

\[e^{i (a+bi)}= \cos(a+bi)+i \sin(a+bi)\] so then we would use sum identities I suppose this might turn ugly

OpenStudy (anonymous):

Okay, now I get: \[ \cos(\Re(z))\cosh(\Im(z))-\cos(\Re(z))\sinh(\Im(z)) \]

OpenStudy (watchmath):

\(e^{-\mathcal{I}(z)}\cdot \cos(\mathcal{R}(z))\)

OpenStudy (kainui):

Nope to all. =P There is only one real and imaginary part and they exist like this and this only \[\Large f(z) = \Re(z) + i \Im(z)\]

OpenStudy (freckles):

\[e^{i(a+bi)}=\cos(a)\cos(bi)-\sin(a)\sin(bi)+i(\sin(a)\cos(bi)+\cos(a)\sin(bi)) \\ =\cos(a)\cosh(b)-\sin(a) \cdot i \sin(x)+i(\sin(a)\cosh(b)+\cos(a) \cdot i(\sin(b)) \\ =\cos(a)\cosh(b)-\sin(b)\cos(a)-\sin(a)\sin(x)i+\sin(A)\cosh(b)i\]

OpenStudy (kainui):

@freckles is on the right path, but he's forgotten exponent rules. =P

OpenStudy (freckles):

so the real part is cos(a)cosh(b)-sin(b)cos(a)?

OpenStudy (freckles):

if i didn't make a mistake

OpenStudy (freckles):

oh

ganeshie8 (ganeshie8):

freckles is she

OpenStudy (freckles):

freckles can be a he who knows

OpenStudy (freckles):

and yes I totally forgot the exponent thingy

OpenStudy (anonymous):

This equation is not true\[\Large f(z) = \Re(z) + i \Im(z)\]

OpenStudy (kainui):

@freckles ,whatever this unicorn may be, you might like that you can derive trig identities this way!\[\Large e^{i(\theta + \phi)} = \cos(\theta + \phi) +i \sin(\theta + \phi) \\ \Large e^{i \theta}e^{i \phi}=(\cos \theta +i \sin \theta) (\cos \phi + i \sin \phi)\]

OpenStudy (kainui):

@wio lol unfortunately you're wrong.

OpenStudy (kainui):

Every complex function has a real and imaginary part.

OpenStudy (anonymous):

Let\(f(z) = z^2\). Then \(f(i) = -1\), and not \(\Re(i) + \Im(i)i = (0)+(1)i = i\).

OpenStudy (anonymous):

I'm fairly certain that: \[ \Re(e^z)=\cos(\Re(z))\cosh(\Im(z))-\cos(\Re(z))\sinh(\Im(z)) \]

OpenStudy (kainui):

One way to find an explicit formula for Re and Im is to just realize that the conjugate negates the imaginary part. \[\Large \bar z = a-bi\]\[\Large z + \bar z = 2a \\...so \\ \Large \Re(z) = \frac{z+ \bar z}{2}\] Similarly we have \[\Large \Im (z) = \frac{z - \bar z}{i2}\]

OpenStudy (freckles):

\[e^{i(a+bi)}=e^{ai+bi^2}=e^{-b}e^{ai}=e^{-b}(\cos(a)+i \sin(a)) \\ e^{-b} \cos(a)+i e^{-b} \sin(a)\] this seems to easy did i fail

OpenStudy (freckles):

the real part is e^(-b)cos(a)

OpenStudy (kainui):

Yes perfect!

OpenStudy (watchmath):

which is what I wrote above

OpenStudy (anonymous):

And I wrote it before watchmath wrote it, but it is wrong, lol

OpenStudy (freckles):

i don't know what the fancy stuff is :p

OpenStudy (kainui):

I was asking for Re(z), if you give me an answer in terms of R(z) then it isn't the answer.

OpenStudy (anonymous):

No, you were asking for \(\Re(f(z))\) and you are probably going to need \(\Re(z)\) and \(\Im(z)\).

OpenStudy (watchmath):

what do you mean? it is the same a = imaginary part of z= I(z) and b=real part of z =R(z) so \(e^{-I(z)}\cos(Re(z))\) is the same

OpenStudy (anonymous):

Remember that \(a=\Re(z)\) and \(b=\Im(z)\)

OpenStudy (kainui):

Yeah you're right, I guess I'll just have to give you all a third of a medal.

OpenStudy (kainui):

However since 1/3 rounds down to 0, I guess nobody gets medals. =(

OpenStudy (anonymous):

But @Kainui , I'm not sure that @freckles and @watchmath . In fact, I think this is a part where Euler's equation somehow breaks down.

OpenStudy (anonymous):

Okay, actually, let me explain a bit better

OpenStudy (anonymous):

What @freckles gave was:\[ f(z) = e^{iz} \]

OpenStudy (kainui):

Oh? Keep going.

OpenStudy (anonymous):

There are two ways to go about doing this problem, one is use exponent algebra, and the other is to use trig sum formula and hyperbolic trig functions

OpenStudy (anonymous):

They should lead to the same result, but I want to try it out on both to be sure.

OpenStudy (watchmath):

ok, I just gave wio a medal :D

OpenStudy (kainui):

Oh wio by the way I just discovered the derivation of the Cauchy-Riemann equations, it's almost obvious after you see it lol.

OpenStudy (anonymous):

Here is one way to do it:\[ e^{z} = e^{i((a+ib)/i)} = e^{i(b-ai)} = e^{bi+a}=e^{a}e^{ib} = e^a\cos(b)+e^ai\sin(b) \]

OpenStudy (anonymous):

Wait, wtf did I just do, that was sort of a waste.

OpenStudy (kainui):

Just take the partial derivative\[\Large \frac{\partial f}{\partial x}=\frac{\partial f}{\partial z}\frac{\partial z}{\partial x} = \frac{\partial f}{\partial z} \\ \Large \frac{\partial f}{\partial y}=\frac{\partial f}{\partial z}\frac{\partial z}{\partial y} = \frac{\partial f}{\partial z} i\] Then set the right hand sides equal after dividing the second one by i to get: \[\Large \frac{ \partial f}{\partial x} = \frac{1}{i} \frac{\partial f}{\partial y}\] and there are your two equations after you plug in f=u+iv.

OpenStudy (anonymous):

Lol, wtf? Okay, but then:\[ e^z = e^{i(z/i)} = \cos(z/i)+i\sin(z/i) \]

OpenStudy (kainui):

I think it is kind of nice, then we get relations like this: \[\Large \cos(ix)=\cosh(x)\] So you can think of how a circle is perpendicular to a hyperbola in a conic section maybe to make sense of this?

OpenStudy (anonymous):

\[ \cos(b-ai)+i\sin(b-ai) = \cos(b)\cosh(a)+\ldots \]Don't you think it is weird how we didn't have \(\cosh\) in our previous solution?

OpenStudy (anonymous):

@Kainui There is something fishy going on, don't you think?

OpenStudy (anonymous):

Do you agree with this?\[ \Re(e^{a+bi}) = e^a\cos(b) \]

OpenStudy (anonymous):

It certainly makes sense.

OpenStudy (kainui):

No, that's what I just said right before that. Just think of it as being on conic sections.

OpenStudy (anonymous):

Wait, so they're not equal?

OpenStudy (kainui):

Huh, no they are equal. \[\Large \cos(ix)=\cosh(x)\]

OpenStudy (anonymous):

Okay, well hold on, I want to work this out.

OpenStudy (anonymous):

First, it is tempting to say: \[ e^z=e^{a+bi} = e^a\bigg(\cos(b)+i\sin(b)\bigg) \]But consider: \[ e^{z} = e^{i(z/i)} = \bigg(\cos(z/i)+i\sin(z/i)\bigg) = \bigg(\cos(b-ai)+i\sin(b-ai)\bigg) \]Will they result in same answer? Maybe so.

OpenStudy (kainui):

I don't know I am just making this up as I go, but it seems like we can figure it out from this picture: